***Please type answer if possible***

GAS VOLUME STOICHIOMETRY PROBLEMS

1. The combustion of ethanol (C2H5OH) takes place by the following reaction equation.

            C2H5OH (l)    +     3 O2 (g)     →    2 CO2 (g)    +    3 H2O (g)

What is the volume of CO2 gas produced by the combustion of excess ethanol by 23.3 grams of O2gas at 25oC and 1.25 atm?

GAS VOLUME STOICHIOMETRY PROBLEMS

2. Acetic acid (CH3COOH) is formed from its elements by the following reaction equation:

            C (graphite)    +     2 H2 (g)     +     O2 (g)     →     CH3COOH (g)

What is the volume of acetic acid CH3COOH gas produced by the reaction of 18.6 grams of H2 gas at 35oC and 1.05 atm?

GAS VOLUME STOICHIOMETRY PROBLEMS

3. The formation of hydrazine (N2H4) from its elements takes place by the following reaction equation.

                        N2 (g)              +                      2 H2 (g)           →                    N2H4 (g)

What are the volumes of N2 gas and H2 gas required to form 28.5 grams of N2H4 at 30oC and 1.50 atm?

GAS VOLUME STOICHIOMETRY PROBLEMS

4. The combustion of benzene (C6H6) takes place by the following reaction equation.

            2 C6H6 (g)     +     15 O2 (g)     →     12 CO2 (g)     +     6 H2O (g)

What is the    volume of CO2 gas formed by the combustion of 18.5 grams of C6H6 at 30oC and 1.50 atm?

PARTIAL PRESSURE - MOLE FRACTION PROBLEMS

5. A mixture of gases consists of 4.00 moles of He, 2.00 moles of H2, 3.00 moles of CO2 and 5.00 moles of Ar. The total pressure of the mixture is 2900 mm. Determine the mole fraction of each gas in the mixture. Determine the mole percent of each gas in the mixture. Determine the partial pressure of each gas in the mixture.

PARTIAL PRESSURE - MOLE FRACTION PROBLEMS

6. Determine the total pressure of a mixture of 0.400 mole of He and 0.600 mole of Ne in a 2.00 liter container at 25oC.

PARTIAL PRESSURE - MOLE FRACTION PROBLEMS

7. Determine the mass of Ar in 1.00 liter of a gas mixture at 25oC which contains 0.300 atm of Ne and has a total pressure of 4.00 atm.

PARTIAL PRESSURE - MOLE FRACTION PROBLEMS

8. A 1.00 liter container holds a mixture of 0.52 mg of He and 2.05 mg of Ne at 25oC. Determine the partial pressures of He and Ne in the flask. What is the total pressure?
Answer #1

since no particular question has been mentioned, so i am answering the first one

1. C2H5OH (l) + 3O2   2CO2 (g) + 3H2O (g)

we are given that mass of oxygen gas = 23.3 g

we know that, number of moles = mass / molar mass ; also molar mass for  O2 = 32.00 g/mol

n(O2 ) = 23.3g / (32 g/mol)

n(O2) = 0.728 mol

from balanced rxn, it can be seen that 3 moles of O2 give 2 moles of CO2

So, moles of CO2  produced by 0.728 moles of O2 = (2/3) X 0.728 mol = 0.485 mol

Volume of CO2  can be calculated from ideal gas equation,

PV = nRT   V = nRT / P

where, n= number of moles, P= pressure = 1.25 atm; V=volume of the gas;

T=Absolute temperature( in K) = (25+273.15) K = 298.15 K

R=Ideal Gas constant also known as Boltzmann Constant = 0.082057 L atm K-1 mol-1.

V = [(0.485 mol) X (0.082057 L atm K-1 mol-1) X (298.15 K)] / 1.25 atm

V = 9.492 L

Question 2, 3 and 4 can be solved in a similar manner.

hope this helps!